OpenStudy (anonymous):

Is 2x+5 a factor of x^3+2x^2+5x-5? Explain why or why not.

jimthompson5910 (jim_thompson5910):

If 2x+5 was a factor, then x = -5/2 would be a root This is because 2x+5 = 0 turns into x = -5/2 when you solve for x

jimthompson5910 (jim_thompson5910):

So all you need to do is plug in x = -5/2 and evaluate. If you get 0 as a final result, then 2x+5 is a factor

OpenStudy (anonymous):

got it, i didnt know you could just make 2x+5=0 and then solve, the 2x was a little confusing. thanks.

jimthompson5910 (jim_thompson5910):

yw the reason why you can do it is say q(x) times 2x+5 would give you x^3+2x^2+5x-5 in other words x^3+2x^2+5x-5 = (2x+5)(q(x)) So if x^3+2x^2+5x-5 = 0, then (2x+5)(q(x)) = 0 2x+5 = 0 or q(x) = 0 etc etc

OpenStudy (anonymous):

ah got it

OpenStudy (agent0smith):

You can also check it by synthetic division, if you've done that.

OpenStudy (anonymous):

my final answers was -20.625, after plugging in -5/2

jimthompson5910 (jim_thompson5910):

so that means that 2x+5 isn't a factor

OpenStudy (anonymous):

yeah